1. (future value) Calculate the future value of $870 invested at 13.5% for two years.

2. (future value) you plan to buy a camping trailer and con afford to set aside $1,320 toward the purchase today. If the annual interest rate is 10.1% compounded every week, how much can you send in three years on the purchase?
Answer #1

1.

Present Value = PV = $870

Interest Rate = r = 13.5% = 0.135

Number of Years = n = 2

Hence, Future Value = FV = PV(1+r)n = 870(1+0.135)2 = $1120.76

2.

Present Value = PV = $1320

Number of weeks in an Year = 52

Number of periods = n = 3*52 = 156

Weekly Interest Rate = r = 10.1%/52 = 0.101/52

Hence, Future Value = FV = PV(1+r)n = 1320(1+0.101/52)156 = $1786.64